There are 52 cards in a standard deck of playing cards. Half are black and half are red. What is the probability of choosing a red card first, keeping it, and then choosing a black card?
Answer by Alan3354(69443)

There are 52 cards in a standard deck of playing cards. Half are black and half are red.
---
Is that right? Thx for the info.
---
What is the probability of choosing a red card first,
26/52 = 1/2
============
keeping it, and then choosing a black card?
26/51
--> (1/2)*(26/51) = 13/51